OpenStudy (germandragon):

The area of a rectangle is 1.7 square units. The length is 3.4 units and the width is z units. Which statements are true? Select each correct answer. A. z=0.5 B. 3.4z=1.7 C. 1.7z=3.4 D. z=5.78

OpenStudy (chrishogan):

Welcome to OpenStudy!!! I can assist you in your question. :)

OpenStudy (germandragon):

yeah please!

OpenStudy (chrishogan):

Is there multiple answers that can be picked? Because two of them are correct.

OpenStudy (germandragon):

yes! I'm better with the all number questions but this one got me. ;-;

OpenStudy (chrishogan):

A = 1.7=3.4*z C= z=0.5

OpenStudy (germandragon):

omg o0o how did you get that?

OpenStudy (chrishogan):

Opps

OpenStudy (chrishogan):

Hold on my letters were wrong.

OpenStudy (chrishogan):

It's A and B. :)

OpenStudy (germandragon):

omg o0o how did you get that?

OpenStudy (chrishogan):

"The area of a rectangle is 1.7 square units. The length is 3.4 units and the width is z units." Your area is 1.7 units. Length = 3.4 and Width = z. To find area the formula is \[a=l*w\]. So if a=1.7, L=3.4 and w=z Your equation would look like this.\[1.7=3.4w\]

OpenStudy (germandragon):

oh I understand

OpenStudy (chrishogan):

But we're not worried about what the area is for the type of equation, we're looking for what the variable Z is equal to. So how would we separate 3.4 from z?

OpenStudy (chrishogan):

\[\frac{ 1.7 }{ 3.4 }=\frac{ 3.4 }{ 3.4 }z\] This should give you...\[0.5=z\]